OpenStudy (anonymous):

sqrt(399), how do i solve it using linear approximation?

OpenStudy (anonymous):

Can you think of a number close to 399 whose square root would be easy to calculate?

OpenStudy (anonymous):

400

OpenStudy (anonymous):

Do you know the 'formula' that you could use? We will use the fact that sqrt(400)=20

OpenStudy (anonymous):

Does \[y - y_{0} = f'(x_{0}) \times (x - x_0)\] look familiar?

OpenStudy (anonymous):

Our \[f(x) = x^{\frac{1}{2}}\] here, as we are trying to approximate a square root. So we could do \[f(399) \approx f(400) + f'(400) \times (399-400)\] Do you see why?